# Unsourced Random Access: A Recent Paradigm for Massive Connectivity

### v1

@@ IV-B Existing Solutions for GMAC URA
As indicated in the previous subsection, it is shown in [36] that conventional random access techniques are energy inefficient for the URA setup. As a result, there have been substantial efforts to develop low-complexity and energy-efficient schemes that can be used in practical systems.
@@
One of the design strategies is slotting the transmission frame, where each active user randomly picks one or a few slots to transmit its data depending on the exact setting. In this direction, Ordentlich and Polyanskiy propose the first practical scheme for URA in [94], where a concatenated coding scheme is considered combined with the $T$-fold ALOHA protocol. $T$-fold ALOHA is an approach similar to slotted ALOHA in the sense that the users choose a random sub-block (slot) to transmit their packets, but in the slotted ALOHA, packets are lost in the case of a collision. However, in $T$-fold ALOHA, $T$ or fewer users can be simultaneously decoded in each slot. In [94], each user randomly chooses one slot to transmit its message. The concatenated coding scheme consists of an inner binary linear code which is used to decode the modulo-2 sum of the codewords transmitted within a slot and an outer code to recover the individual messages. Many off-the-shelf codes can be used for the inner code. The authors construct the outer code from the columns of a $T$-error correcting Bose–Chaudhuri–Hocquenghem (BCH) code. This scheme suffers in terms of energy efficiency as the performance gap with the random coding bound is about 20 dB when the number of active users is 300. Its performance is significantly improved in [95], where the authors employ a concatenated coding scheme with polar and BCH codes as the inner and outer codes, respectively, and propose a decoding approach exchanging the information between inner and outer decoding.
@@
The authors of [64] consider a slotted structure where the users can repeat their codewords across sub-blocks, which is referred to as T-fold IRSA. They propose to split the message into two parts and employ a coding scheme based on a combination of compressed sensing (CS) and LDPC coding at the slot level. Namely, the users pick a preamble determined by the first part of the message from a common codebook, and the second part is encoded by an LDPC code and interleaved before the transmission. The preamble carries information about the interleaver and the utilized slots for a specific user. At the receiver side, the preamble part is recovered by the non-negative least squares (NNLS) algorithm and the coded part is decoded by joint belief propagation (BP). The replicas of the recovered codewords are peeled from the other slots via SIC. The general transmission structure in CCS-based URA solutions is depicted in Fig. 6.
@@
Another main approach in URA over GMAC is called coded compressed sensing (CCS) [55], [97, 98, 99, 100, 101, 102, 103], where the messages are divided into sub-blocks and each block is first encoded by an outer code by appending redundant parity bits before being mapped to a CS codebook for the inner encoding. The encoded sub-blocks are then transmitted in different slots. At the receiver side, each sub-block is recovered via compressed sensing algorithms, and the recovered segments are stitched together by an outer tree code. Based on this main idea, several variations are examined in the literature. Specifically, in [55], the sub-blocks are first encoded by a binary linear code and then mapped to the columns of a common sensing matrix for inner encoding. In [98], sparse regression codes (SPARCs) are exploited as inner codes with approximate message passing (AMP) decoding at the receiver. The authors of [99] improve the performance of the proposed scheme in [98] by introducing a dynamic interaction between the inner AMP and the outer tree decoder. SPARCs are also employed in [100], where the sub-blocks are CRC-encoded and connected via block Markov superposition transmission. At the receiver, a hybrid decoder combining SIC and AMP is employed for inner decoding, with the addition of tree decoding. Furthermore, the authors of [101] adapt constant-weight codes as the outer code with the use of Gabor frames to design the inner CS coding matrix, and a concatenation of Raptor codes and tree codes is employed as the outer code in [102]. Finally, in a recent work [103], the authors propose a CCS-based scheme eliminating the tree code by using the encoded sub-message in the first slot as a user identifier. This reduces the parity bits and the resulting rate loss, offering the best performance among the CCS-based schemes on GMAC and outperforming the remaining solutions for $K_{a}\leq 225$. The general encoding-decoding structure in coded compressed sensing is illustrated in Fig. 7.
Figure 7: Encoding-decoding structure in coded compressed sensing (CCS).
Figure 8: A typical decoding structure in random spreading.
@@
In another design strategy for URA over GMAC, there is no slotting and the active users utilize the whole transmission frame. In some works following this strategy, the idea is to spread the channel-coded bits through the transmission frame using random signatures [68, 69]. The message is divided into two parts: the first part determines the signature sequence while the second part is encoded by a polar code in [68]. At the receiver, first, the selected signature sequences are detected by a correlation-based energy detector. Then, an iterative decoding algorithm consisting of MMSE filtering followed by single-user polar decoding is employed to recover the channel-coded bits. Moreover, SIC is applied at the end of each iteration. The general decoding process of random spreading based solutions is illustrated in Fig. 8.
@@
In [104], the authors modify the scheme in [68] by replacing the polar codes with polarization-adjusted convolutional (PAC) codes, with a slight performance improvement and reduced implementation complexity. Moreover, [81] improves the performance of the scheme in [68] by letting the users to randomly select a power level from a pre-defined set. This approach, called power diversity, creates an artificial fading-like scenario, where the users with higher received power can be decoded first, and the ones with lower received power are decoded after SIC. This results in a performance improvement over [68] for $K_{a}\geq 150$. The authors of [69] replace the polar codes in [68] with LDPC codes to exploit the soft information produced by the belief propagation decoder to perform soft-input-soft-output (SISO) MMSE filtering. This approach improves the decoding performance since the information about the partially decoded codewords can be utilized, and it provides the best performance among the schemes based on random spreading.
@@
In random spreading based solutions, each active user occupies the entire frame by random spreading of channel-coded bits for data transmission. In some other works, the active users distribute their codeword bits onto the transmission frame in different ways. For instance, in the sparse interleave-division multiple access (IDMA) approach [67], the message bits and the channel frame are divided into parts. The users transmit a preamble sequence dictated by the first part of their message in the preamble part of the frame. The rest of the bits are encoded by an LDPC code, zero-padded, and permuted by an interleaver determined by the preamble to be transmitted along the rest of the channel frame. This encoding process is illustrated in Fig. 9. At the receiver, after decoding the preamble part via the least absolute shrinkage and selection operator (LASSO) technique, a single joint Tanner graph is utilized for message recovery, exploiting the information of the location of transmitted symbols from the identified preamble.
@@
A similar approach to [67] employing polar codes is adopted in [105], however, different code lengths are used among the users and TIN is utilized as the decoding algorithm. Moreover, in [106], the message bits are first encoded by a high-rate error correction code. The encoded bits are divided into two parts with the first part being mapped to a column of a common sensing matrix which determines the interleaver and signature to be utilized in the second part after its repetition. The receiver utilizes AMP to recover the first part followed by a multiuser decoding algorithm proposed in the second part. Furthermore, the authors of [70] propose to encode the data as the Kronecker product of two-component codewords, and employ an iterative decoder based on bilinear generalized AMP to decompose the Kronecker product and a soft-in-soft-out decoder for the individual components. Furthermore, enhanced spread spectrum ALOHA (E-SSA) [107], an asynchronous random access protocol combining ALOHA with direct sequence spread spectrum is adapted for URA over GMAC in [108]. The authors employ a wrap-around version of E-SSA to be able to compare their proposed solution with the existing synchronous solutions and benchmarks.
In a recent work [109], the authors employ on-off division multiple access (ODMA) [110] in the context of URA, where the active users exploit a small portion of the transmission frame to transmit their data encoded by a repeat-accumulate (RA) code according to a randomly selected transmission pattern, with the remaining time instances being idle. At the receiver, the BS first recovers the transmission patterns and then employs multiuser detection and decoding over a sparse factor graph. The authors of [111] investigate the ODMA technique in URA further, namely, they employ polar codes to improve the performance in the low active user load regimes combined with a simpler pattern detection algorithm. Another high-performing solution is proposed in [112], which improves the scheme in [109] by employing a non-binary LDPC code and resolving the collisions of two users.
A summary of the available coding schemes in URA over GMAC can be found in Table II, accompanied by a comparison of their performance in terms of energy efficiency in Fig. 10 for $n=30000$, $B=100$, and $\epsilon=0.05$, which are the common set of parameters in the literature. The results in Fig. 10 demonstrate that random spreading based schemes (Pradhan et al., 2020, Ahmadi et al. 2021, Pradhan et al., 2021) offer good performance for a lower number of active users; however, their performance degrades with the increasing number of active users since each active user occupies the entire transmission frame. On the other hand, ODMA (Yan et al., 2023, Ozates et al., 2024, Yan et al., 2024) outperforms the other approaches except the solution proposed in Nassaji et al., 2024 in the high multiuser interference regime due to its sparsity, while it performs well in the low multiuser regime when combined with polar coding. Moreover, CCS can exhibit very good performance for a high number of active users when the rate loss due to the redundant parity bits is alleviated as shown in Nassaji et al., 2024. Regarding the comparison with the achievability bound, the proposed schemes can perform similar or slightly better for $K_{a}\leq 100$, and the gap is a maximum of 0.8 dB for $K_{a}\leq 300$.
Table II: Main coding schemes in URA over GMAC
Paper Contribution and Techniques
@@
Polyanskiy, 2017 [36] information-theoretic formulation of URA. Random coding achievability and converse bounds.
@@
Ordentlich et al., 2017 [94] slotting based scheme with concatenated coding combined with $T$-fold ALOHA
Vem et al., 2019 [64] slotting based scheme with LDPC coding, preamble transmission, and SIC
Marshakov et al., 2019 [71] slotting based scheme with preamble transmission, polar coding, and SIC
Amalladinne et al., 2020 [55] CCS-based scheme with inner CS and outer tree coding.
Amalladinne et al., 2022 [99] CCS-based scheme with the interaction between the inner AMP and the outer BP
Nassaji et al., 2024 [103] CCS-based scheme eliminating the outer tree code.
Pradhan et al., 2022 [67] sparse spreading based scheme with preamble transmission, LDPC coding and SIC
Han et al., 2021 [70] sparse spreading based scheme with convolutional coding and SIC
Pradhan et al., 2020 [68] random spreading based scheme with polar coding and SIC
Ahmadi et al., 2020 [81] random spreading based scheme with power diversity, polar coding, and SIC
Pradhan et al., 2021 [69] random spreading based scheme with LDPC coding, soft cancellation, and SIC
Yan et al., 2023 [109] ODMA based scheme with RA coding and soft cancellation
Ozates et al., 2024 [111] ODMA based scheme with polar coding and SIC
Yan et al., 2024 [112] ODMA based scheme with non-binary LDPC coding and soft cancellation.
@@ V URA over Fading Channels
Since GMAC is not a very realistic channel model for wireless communications, a number of studies explore URA over Rayleigh fading channels [67, 83, 84, 114, 116, 115, 113, 56, 72, 74, 75, 73]. In [83, 84, 113, 114, 115], the authors derive some performance bounds for the URA setup over a Rayleigh fading channel. For instance, the authors of [115] present a converse bound based on the results from [117] and meta-converse from [118], stated below.
@@
where $G=\norm{h}^{2}$ and $Z_{i}\sim\mathcal{CN}(0,1)$. Then, for every $n$ and $10<\epsilon<$ 1, any $(M,n-1,\epsilon)$ code for the quasi-static $K_{a}$ multiple access satisfies
@@
where $\gamma_{n}$ is the solution of $\mathbb{P}\left[S_{n}\leq n\gamma_{n}\right]=\epsilon$.                                             $\blacksquare$
@@
In [84] and [113], achievability and converse bounds are proposed for the cases of known and unknown CSI in an asymptotic setting, where the number of users grows linearly with the blocklength. Authors in [114, 115] provide both asymptotic and non-asymptotic achievability and converse bounds for the case of unknown CSI. In [83], authors derive an achievability bound for the URA scheme over the Rayleigh fading channel by considering slotted structure and perfect SIC in their analysis. Their proposed non-asymptotic achievability bound is tighter than the ones in [84, 113, 114, 115] however, the assumption of perfect SIC may not be realistic.
V-A Single-Antenna Fading Channels

### v2

@@ IV-B URA Solutions for GMAC
As shown in the previous subsection, the conventional RA techniques are energy-inefficient for the URA setup. As a result, there have been substantial efforts to develop low-complexity and energy-efficient schemes that can be used in practical systems. The main approaches for this purpose in the literature can be categorized as slotting the transmission frame, coded compressed sensing (CCS), random spreading, sparse spreading, and on-off division MA (ODMA). In the following, we review the existing practical coding schemes for URA over GMAC based on this classification.
IV-B1 Slotting-based schemes
@@
One of the design strategies is slotting the transmission frame, where each active user randomly picks one or a few slots to transmit its data depending on the exact setting. In this direction, Ordentlich and Polyanskiy propose the first practical scheme for URA in [53], where a concatenated coding scheme is considered combined with the $T$-fold ALOHA protocol. $T$-fold ALOHA is an approach similar to slotted ALOHA in the sense that the users choose a random sub-block (slot) to transmit their packets, but in the slotted ALOHA, packets are lost in the case of a collision. However, in $T$-fold ALOHA, $T$ or fewer users can be simultaneously decoded in each slot. In [53], each user randomly chooses one slot to transmit its message. This concatenated coding scheme consists of an inner binary linear code which is used to decode the modulo-2 sum of the codewords transmitted within a slot and an outer code to recover the individual messages. Many off-the-shelf codes can be used for the inner code. The authors construct the outer code from the columns of a $T$-error correcting Bose–Chaudhuri–Hocquenghem (BCH) code. This scheme suffers in terms of energy efficiency as the performance gap with the random coding bound is about 20 dB when the number of active users is 300. Its performance is significantly improved in [54], where the authors employ a concatenated coding scheme with polar and BCH codes as the inner and outer codes, respectively, and propose a decoding approach exchanging the information between inner and outer decoding.
@@
The authors of [55] consider a slotted structure where the users can repeat their codewords across sub-blocks, which is referred to as T-fold IRSA. They propose to split the message into two parts and employ a coding scheme based on a combination of compressed sensing (CS) [56] and LDPC coding at the slot level. Namely, the users pick a preamble determined by the first part of the message from a common codebook, and the second part is encoded by an LDPC code and interleaved before the transmission. The preamble carries information about the interleaver and the utilized slots for a specific user. At the receiver side, the preamble part is recovered by the non-negative least squares (NNLS) algorithm and the coded part is decoded by joint belief propagation (BP). The replicas of the recovered codewords are peeled from the other slots via SIC. The general transmission structure in T-fold IRSA based URA solutions is depicted in Fig. 5.
@@
IV-B2 CCS-based schemes
Figure 6: Encoding-decoding structure in a coded compressed sensing (CCS) approach.
Figure 7: A typical decoding structure in a random spreading approach.
Another main approach in URA over GMAC is called coded compressed sensing (CCS) [60, 61, 62, 63, 64, 65, 66], where the messages are divided into sub-blocks and each sub-block is first encoded by an outer code by appending redundant parity bits before being mapped to a CS codebook for the inner encoding. The encoded sub-blocks are then transmitted in different slots. At the receiver side, each sub-block is recovered via compressed sensing algorithms, and the recovered segments are stitched together by an outer tree code. The general encoding-decoding structure in coded compressed sensing is illustrated in Fig. 6. More specifically, the encoding process is performed in two phases: 1) In the tree encoder, $B$ bits of information are divided into $S$ parts, and all of these segments, except the initial one, are appended by additional parity-check bits; 2) in the CS encoder, each of $S$ sub-message is mapped to a CS codebook to select a sequence for transmission. Hence, the received signal is comprised of $S$ slots, with each slot being a superposition of the sequences from $K_{a}$ users. The decoder also functions in two stages: 1) In the initial stage, $K_{a}$ sequences associated with the $K_{a}$ active users are identified within each slot. 2) in the second stage, called outer decoder or tree decoder, with the assistance of the unique parity bits, the sub-messages corresponding to the detected sequences are combined to create a message of length $B$.
@@ IV-B2 CCS-based schemes
Several variations of the CCS approach are examined in the URA literature. Specifically, in [60], the sub-blocks are first encoded by a binary linear code and then mapped to the columns of a common sensing matrix for inner encoding. In [61], sparse regression codes (SPARCs) are exploited as inner codes with approximate message passing (AMP) decoding at the receiver. The authors of [62] improve the performance of the proposed scheme in [61] by introducing a dynamic interaction between the inner AMP and the outer tree decoder. SPARCs are also employed in [63], where the sub-blocks are CRC-encoded and connected via block Markov superposition transmission. At the receiver, a hybrid decoder combining SIC and AMP is employed for inner decoding, with the addition of tree decoding. Furthermore, the authors of [64] adapt constant-weight codes as the outer code with the use of Gabor frames to design the inner CS coding matrix, and a concatenation of Raptor codes and tree codes is employed as the outer code in [65]. Finally, in a recent work [66], the authors propose a CCS-based scheme eliminating the tree code by using the encoded sub-message in the first slot as a user identifier. This reduces the parity bits and the resulting rate loss, offering the best performance among the CCS-based schemes on GMAC and outperforming the remaining solutions for $K_{a}\geq 225$.
IV-B3 Random spreading-based schemes
@@
In another design strategy for URA over GMAC, there is no slotting and the active users utilize the whole transmission frame. In some works following this strategy, the idea is to spread the channel-coded bits through the transmission time frame using random signatures. The random spreading idea was first introduced in [67], where the message is divided into two parts: the first part determines the signature sequence while the second part is encoded by a polar code. At the receiver, first, the selected signature sequences are detected by a correlation-based energy detector. Then, an iterative decoding algorithm consisting of MMSE filtering followed by single-user polar decoding is employed to recover the channel-coded bits. Moreover, SIC is applied at the end of each iteration. The general decoding process of random spreading-based solutions is illustrated in Fig. 7.
@@
In [68], the authors modify the scheme in [67] by replacing the polar codes with polarization-adjusted convolutional (PAC) codes, with a slight performance improvement and reduced implementation complexity. Moreover, [44] improves the performance of the scheme in [67] by letting the users to randomly select a power level from a pre-defined set. This approach, called power diversity, creates an artificial fading-like scenario, where the users with higher power levels are decoded first with a high probability, and the ones with lower power levels are decoded next after SIC. This results in a performance improvement over [67] for $K_{a}\geq 150$. The authors of [69] replace the polar codes in [67] with LDPC codes to exploit the soft information produced by the belief propagation decoder to perform soft-input-soft-output (SISO) MMSE filtering. This approach improves the decoding performance since the information about the partially decoded codewords can be utilized, and it provides the best performance among the schemes based on random spreading. Furthermore, enhanced spread spectrum ALOHA (E-SSA) [70], an asynchronous RA protocol combining ALOHA with direct sequence spread spectrum is adapted for URA over GMAC in [71]. The authors employ a wrap-around version of E-SSA in order to adapt it for a synchronous URA setup.
IV-B4 IDMA-based schemes
@@
In random spreading based solutions, each active user occupies the entire time frame by random spreading of channel-coded bits for data transmission. In some other works, the active users distribute their codeword bits onto the transmission time frame in different ways. For instance, in the sparse interleave-division MA (IDMA) approach [72], the message bits and the time frame are divided into parts. The users transmit a preamble sequence dictated by the first part of their message in the preamble part of the frame. The rest of the bits are encoded by an LDPC code, zero-padded, and permuted by an interleaver determined by the preamble to be transmitted along the rest of the channel frame. This encoding process is illustrated in Fig. 8. At the receiver, after decoding the preamble part via the least absolute shrinkage and selection operator (LASSO) technique, a single joint Tanner graph is utilized for message recovery, exploiting the information of the location of transmitted symbols from the identified preamble.
@@
A similar approach to the sparse IDMA in [72] is adopted in [73] employing polar codes, however, different code lengths are used among the users and TIN is utilized as the decoding algorithm. Moreover, in [74], the message bits are first encoded by a high-rate error correction code. The encoded bits are divided into two parts, with the first part being mapped to a column of a common sensing matrix, which determines the interleaver and signature to be utilized in the second part after its repetition. The receiver utilizes AMP to recover the first part followed by a multiuser decoding algorithm for the second part. Furthermore, the authors of [75] propose to encode the data as the Kronecker product of two-component codewords, and employ an iterative decoder based on bilinear generalized AMP to decompose the Kronecker product and a soft-in-soft-out decoder for the individual components.
Table III: Main URA solutions for GMAC
Paper Approach Contribution and Techniques
@@
Polyanskiy, 2017 [11] - information-theoretic formulation of URA, and Random coding achievability and converse bounds
@@
Ordentlich et al., 2017 [53] slotting concatenated coding combined with $T$-fold ALOHA
Vem et al., 2019 [55] slotting LDPC coding, preamble transmission, and SIC
Marshakov et al., 2019 [57] slotting preamble transmission, polar coding, and SIC
Amalladinne et al., 2020 [60] CCS inner CS and outer tree coding
Amalladinne et al., 2022 [62] CCS interaction between the inner AMP and the outer BP
Nassaji et al., 2024 [66] CCS eliminating the outer tree code
Pradhan et al., 2022 [72] sparse spreading preamble transmission, LDPC coding and SIC
Han et al., 2021 [75] sparse spreading convolutional coding and SIC
Pradhan et al., 2020 [67] random spreading polar coding and SIC
Ahmadi et al., 2020 [44] random spreading power diversity, polar coding, and SIC
Pradhan et al., 2021 [69] random spreading LDPC coding, soft cancellation, and SIC
Yan et al., 2023 [76] ODMA repeat-accumulate coding and soft cancellation
Ozates et al., 2024 [78] ODMA polar coding and SIC
Yan et al., 2024 [79] ODMA non-binary LDPC coding and soft cancellation
IV-B5 ODMA-based schemes
Another way to distribute the user codeword bits onto the transmission time frame is proposed in [76], where the authors employ on-off division MA (ODMA) [77] in the context of URA. In this approach, the active users exploit a small disjoint portion of the transmission time frame to transmit their data encoded by a repeat-accumulate code according to a randomly selected transmission pattern, with the remaining time instances being idle. At the receiver, the BS first recovers the transmission patterns and then employs multiuser detection and decoding over a sparse factor graph. The authors of [78] investigate the ODMA technique in URA further, namely, they employ polar codes to improve the performance in the low active user load regimes combined with a simpler pattern detection algorithm. Another high-performing solution is proposed in [79], which improves the scheme in [76] by employing a non-binary LDPC code.
@@ IV-B5 ODMA-based schemes
IV-C Summary and Performance Comparison
A summary of main URA solutions for GMAC can be found in Table III, accompanied by a comparison of their performance in terms of energy efficiency in Fig. 9 for $n=30000$, $B=100$, and $\epsilon=0.05$, which are the common set of parameters in the literature. The results in Fig. 9 demonstrate that random spreading-based schemes in [67, 44, 69] offer good performance for a lower number of active users; however, their performance degrades with the increasing number of active users since each active user occupies the entire transmission frame. On the other hand, ODMA-based approaches in [76, 78, 79] outperform the others except the solution proposed in [66] in the high multiuser interference regime due to its sparsity, while it performs well in the low multiuser regime when combined with polar coding. Moreover, CCS-based approaches can exhibit very good performance for a high number of active users when the rate loss due to the redundant parity bits is alleviated, as shown in [66]. Compared to the achievability bound, the proposed schemes can perform similar or slightly better for $K_{a}\leq 100$, and the gap is a maximum of 0.8 dB for $K_{a}\leq 300$.
IV-D Computational Complexity
In terms of complexity, slotting is the most beneficial approach among the existing solutions for GMAC URA, as the complexity of the operations depending on the user count super-linearly increases by the number of active users in the slot rather than $K_{a}$. In addition, slotting leads to smaller preamble/pilot length and channel code length $n_{c}$, decreasing the complexity of the operations depending on these parameters. CCS also has a low complexity since slot-wise compressed sensing and a low-complexity tree code are employed in conjunction. Namely, its complexity order is linear or quadratic in $K_{a}$ depending on the regime [60]. ODMA is another low-complexity approach, as only a small fraction of the time frame is occupied by each user, which simplifies the iterative multiuser decoding process [76]. More specifically, its complexity order linearly depends on $n_{c}$ and $K_{a}$. On the other hand, random spreading-based schemes where each user occupies the entire time frame have the highest complexity, as the complexity order is cubic in $K_{a}$ due to the MMSE filtering [67]. On another note, the number of computations needed for some decoding operations depends on the transmission time frame length rather than the slot or code length as in slotting or ODMA, which increases the complexity.
@@ V URA over Fading Channels
Since GMAC is not a very realistic channel model for wireless communications, most works in the URA literature consider fading channels, the majority of them dedicated to Rayleigh fading channels. In this section, we first review the bounds on the performance of fading URA solutions that have been provided in the literature. Then, we cover the URA solutions that present low-complexity coding schemes to achieve those bounds. To this end, in order to explain the fading URA approaches in a simpler setup for better understanding, we first cover the single-antenna fading solutions, then, we turn our attention to the MIMO fading ones.
V-A Performance Bounds
In [85, 83, 84, 80], the authors derive some performance bounds for the URA setup over a Rayleigh fading channel. In one of the main analytical results, the authors of [80] present a converse bound for fading URA based on the results in [81] and the meta-converse bound in [82], as stated below.
@@ V-A Performance Bounds
where $G=\norm{h}^{2}$ and $Z_{i}\sim\mathcal{CN}(0,1)$. Then, for every $n$ and $10<\epsilon<$ 1, any $(M,n-1,\epsilon)$ code for the quasi-static $K_{a}$ MA satisfies
@@
where $\gamma_{n}$ is the solution of $\mathbb{P}\left[S_{n}\leq n\gamma_{n}\right]=\epsilon$.
For the proof, first, consider the fact that the converse bound when channel state information (CSI) is available at the receiver is a converse for the case with no CSI case as well. Moreover, by symmetry, it is enough to lower bound the probability that a particular user’s message is not in the decoded list. Finally, assuming that the decoder knows the codewords of all the other users (which makes this in essence a single-user bound), the rest of the proof becomes quite straightforward, which can be found in [80]. The authors numerically show that the $E_{b}/N_{0}$ requirements for a fading channel are necessarily higher compared to the AWGN channel. They also mention that, for the common system parameters in a fading URA setup, a Fano-type converse bound is worse than the one presented in Theorem 2.
@@
In [83] and [84], achievability and converse bounds are proposed for the cases of known and unknown CSI in an asymptotic setting, where the number of users grows linearly with the blocklength. The authors in [80] provide both asymptotic and non-asymptotic achievability and converse bounds for the case of unknown CSI. In [85], authors derive an achievability bound for the URA scheme over the Rayleigh fading channel by considering slotted structure and perfect SIC in their analysis, which is tighter than the other existing bounds. However, since the channel coefficients used in a SIC process are estimated, hence imperfect, the assumption of perfect SIC is not realistic, which makes the result in [85] an approximation result rather than an achievability one.
Moreover, in [86, 87], achievability and converse bounds are obtained for URA over MIMO Rayleigh fading channels, whose ML-based achievability bound is obtained by employing a projection decoder presented in [88].
V-B Single-Antenna Fading Channels